divide 3/4 by 1/2, 1/4, 1/8, and 1/12
Answer by Theo(13342)

dividing by a number is the same as multiplying by its reciprocal.

for example:

5/4 = 5 * 1/4

this works with fractions as well.

5 / (1/4) = 5 * 4 because the reciprocal of 1/4 is 4.

5 / (2/3) = 5 * 3/2 because the reciprocal orf 2/3 is 3/2.

the same concept is applied to your problems.

(3/4) / (1/2) is equal to (3/4) * (2/1) which is equal to (3*2) / (4*1) which is equal to 6/4 which can be simplified to 3/2.

(3/4) / (1/4) is equal to (3/4) * (4/1) which is equal to (3*4) / (4*1) which is equal to 12/4 which is equal to 3.

(3/4) / (1/8) is equal to (3/4) * (8/1) which is equal to (3*8) / (4*1) which is equal to 24 / 4 which is equal to 6.

(3/4) / (1/12) is equal to (3/4) * (12/1) which is equal to (3*12) / (4*1) which is equal to 36 / 4 which is equal to 9.

you can confirm using your calculator.

don't forget to use parentheses when you make your entries into the calculator if you do it all in one shot.

using the calculator:

(3/4) / (1/2) = 1.5 which is the same as 3/2.

(3/4) / (1/4) = 3

(3/4) / (1/8) = 6

(3/4) / (1/12) = 9